OpenStudy (candy13106):

Write an equation of the line that passes through the given point and is perpendicular to the given line. (2,9),4x+y=3

OpenStudy (candy13106):

@MrNood

OpenStudy (candy13106):

@dan815 @esshotwired @Walkerman15 @CloverRacer @Crissy15 @Kit_Kat_Nat_<3

OpenStudy (candy13106):

@Whiterabbit541 @whydoihavetosignup1 @Technogirl123 @mathmath333

OpenStudy (candy13106):

help

OpenStudy (anonymous):

Perpendicular lines have negative reciprocals slopes of each other. So, the slope of the equation of the line perpendicular to y=-2x+7 is 1/2. Now to find the y-intercept, all you have to do is plug in the x and y values of (2,-4). The slope intercept form is y=mx+b, where m is the slope and b is the y-intercept. y=1/2x+b -4=(1/2)2+b -4=2/2+b -4=1+b -5=b Thus, the the equation of the line that passes through the given point and is perpendicular to the given line is y=1/2x-5. y=1/2x-5 y-1/2x=-5 Since, the A has to be positive. Multiply -1 on both sides. y-1/2x=-5 -y+1/2x=5 Thus, in standard form, it is -y+1/2x=5 . Hope I helped! :)

OpenStudy (candy13106):

thanks

OpenStudy (anonymous):

Welcome